Need help with this one and not sure how to work out the problem.

Suppose you travel north for 35 kilometers then travel east 65 kilometers. How far are you from your starting point? North and east can be considered the directions of the y- and x-axis respectively. Round to the tenth's place.

Answer by stanbon(48564)

Suppose you travel north for 35 kilometers then travel east 65 kilometers. How far are you from your starting point? North and east can be considered the directions of the y- and x-axis respectively. Round to the tenth's place.
--------------
Draw the picture.
The distance you are looking for is the hypotenuse of
of a right triangle with sides of length 35 and 65.
EQUATION:
d^2=35^2+65^2
d^2=5450
d=73.82 Km
Cheers,
Stan H.